a) 0.154g of organic compound Containing calbon, Thydrogen and oxygen only, on ignition in oxygen gake 0-2149 of carbon (iv) oxile and 0.135g of Loater. If thes Vapour dansity of the compound is 31 formula (b) it's molecular formula It's emperical


b Calculate the number of grammes of substance present in 3rol + 1022 atoms of letra osco sulphate (vi) acid


1
Expert's answer
2021-11-08T08:23:43-0500

Molar mass of tetraoxosulphate (VI) acid = 142.04 g/mol

3×1022/6.022×1023= 0.0498mol

= 142.04×0.0498= 7.0761g